# Distributed Hierarchical Machine Learning for Joint Resource Allocation and Slice Selection in In-Network Edge Systems

### v1

@@ III System Model
We consider a slice-enabled network edge domain characterized by distinct application slices $\mathcal{N}=\{1,2,\ldots,N\}$, which have combinations of computing resources optimized for executing intensive tasks. We denote by $\mathcal{I}=\{1,2,\ldots,I\}$ as set of Wireless Devices (WD) that generate computationally intensive tasks each with varying computational demands. Task $i$ generated by II characterized with input size $S_{i}$ can either be computed locally, or assigned to a slice $n$ through set $\mathcal{A}=\{1,2,\ldots,A\}$ of access points (AP).
@@ III-C Cost Model
$\begin{split}C(\boldsymbol{\delta},\mathbb{R}_{\omega},\mathbb{R}_{\phi_{a}},\mathbb{R}_{\phi_{j}})=&\sum_{i\in I}C_{i}(\boldsymbol{\delta},\mathbb{R}_{\omega},\mathbb{R}^{n}_{\phi_{a}},\mathbb{R}^{n}_{\phi_{j}})\\ &=\sum_{n\in N}C_{(n)}(\boldsymbol{\delta},\mathbb{R}_{b},\mathbb{R}^{n}_{\phi_{a}},\mathbb{R}^{n}_{\phi_{j}})\\ &+\sum_{i\in O_{l}(\delta)}C_{i}^{l}\end{split}$ (12)
@@ V Distributed Hierarchical DeepSet-S model for resource management and task offloading
Since the WD count of II differs across sets, we pad each $X$ to a common set size $\mathcal{I}_{\max}$ within a mini-batch. Let $\mathrm{pad}(\cdot)$ append all-zero rows as needed; the padded matrix is
@@
Given $(\tilde{X},m)$ for a set with $\mathcal{I}$ real WD $i$ and a fixed set of allocation policy $\mathbb{R}$ indexed by $e\in\{1,\dots,\mathcal{E}\}$, our objective is to learn a permutation-equivariant mapping that outputs, for each $\mathbb{R}$, a probability distribution over II plus a slack (unused capacity) entry, summing to one per $\mathbb{R}$. Formally,
@@ V-B Decoder
The SP1 decoder generates, for each resource $e$ in slice $n$, an augmented softmax distribution over the index set $\{0\}\cup\mathcal{I}$, where index 0 denotes a slack option corresponding to unused capacity. Let the resulting probabilities be $p_{0},p_{1},\ldots,p_{|\mathcal{I}|}$ with $p_{k}\geq 0$ and $\sum_{k}p_{k}=1$. Define $\phi^{n}_{i,e}=p_{i}$ for each user ii. Then,
@@
Let the SP3 classifier output a probability vector $p_{i}=[p_{i}^{1},p_{i}^{2},\ldots,p_{i}^{K}]\in\mathbb{R}^{K}$, where $K=|\Delta_{i}|$ is the number of admissible offloading decisions for WD ii. Since the final decision is taken as $\widehat{\delta}_{i}=\arg\max_{k}p_{i}^{k}$, exactly one index is active while all others are inactive. Therefore,
@@ VI-B Experiment Environment
We consider a small-scale scenario for an edge application with a square area of (500 x 500)$m$ where 8 COINs, 1 MEC, 3 APs, and WDs are placed at random. The power spectral density of Gaussian noise $N_{0}$ is 174dBm/Hz-174\ dBm/Hz [aliyu2024digital], and according to [jovsilo2022joint], we represent the achievable physical rate $R_{i,a}$ from WD $i$ to AP $a$ as $R_{i,a}=B_{a}log\ (1+(distance)P^{t}/N_{0})$, where distancedistance is the Euclidean distance between WD $i$ and AP $a$, and $P^{t}$ is the transmission power of WD ii. Other simulation parameters are summarized in table II.
@@
  • Proportional Resource Allocation Policy (PRAP): We define the policies $\mathbb{R}_{\omega}^{pr}$, which share the bandwidth of each AP $a$ to each slice $n$ proportionally, $\mathbb{R}_{\phi_{a}}^{pr}$ which gives propotional share of the bandwidth of each AP $a$ in slice $n$ to each WD $i$, and finally $\mathbb{R}_{\phi_{j}}^{pr}$ which gives proportional share of the computing resource of node $j$ in slice $n$ to each WD ii.
  • Equal Resource Allocation Policy (ERAP): We define the policies $\mathbb{R}_{\omega}^{eq}$, which gives equal share of the bandwidth of each AP $a$ to each slice $n$, $\mathbb{R}_{\phi_{a}}^{eq}$ which gives equal share of the bandwidth of each AP $a$ in slice $n$ to each WD $i$, and finally $\mathbb{R}_{\phi_{j}}^{eq}$ which gives equal share of the computing resource of node $j$ in slice $n$ to each WD ii.
@@
To evaluate the overall efficiency of our proposed framework, we define the system cost ratio CRCR as a key performance metric. The CRCR quantifies the ratio between the system cost reached using the optimization solver ($C_{OS}$) and the system cost reached under the baseline model ($C_{m}$).
@@
A higher CRCR indicates that the model is more effective at minimizing system costs and is performing closer to the optimal solution.
Figure 6 illustrates the CRCR as a function of the number of APs for WD sizes of 15 and 25. In both scenarios, the proposed DeepSets-S model consistently achieves higher cost ratios than the baselines ERAP and PRAP, confirming its ability to approximate the optimal solution more closely. The performance gap is more pronounced for larger workloads (WD = 25), where DeepSets-S maintains a substantial advantage.

### v2

@@ III System Model
We consider a slice-enabled network edge domain characterized by distinct application slices $\mathcal{N}=\{1,2,\ldots,N\}$, which have combinations of computing resources optimized for executing intensive tasks. We denote by $\mathcal{I}=\{1,2,\ldots,I\}$ as set of Wireless Devices (WD) that generate computationally intensive tasks each with varying computational demands. Task $i$ generated by $I$ characterized with input size $S_{i}$ can either be computed locally, or assigned to a slice $n$ through set $\mathcal{A}=\{1,2,\ldots,A\}$ of access points (AP).
@@ III-C Cost Model
$\begin{split}C(\boldsymbol{\delta},\mathbb{R}_{\omega},\mathbb{R}_{\phi_{a}},\mathbb{R}_{\phi_{j}})=&\sum_{i\in I}C_{i}(\boldsymbol{\delta},\mathbb{R}_{\omega},\mathbb{R}^{n}_{\phi_{a}},\mathbb{R}^{n}_{\phi_{j}})\\ &+\sum_{n\in N}C_{(n)}(\boldsymbol{\delta},\mathbb{R}_{b},\mathbb{R}^{n}_{\phi_{a}},\mathbb{R}^{n}_{\phi_{j}})\end{split}$ (12)
@@ V Distributed Hierarchical DeepSet-S model for resource management and task offloading
Since the WD count of $I$ differs across sets, we pad each $X$ to a common set size $\mathcal{I}_{\max}$ within a mini-batch. Let $\mathrm{pad}(\cdot)$ append all-zero rows as needed; the padded matrix is
@@
Given $(\tilde{X},m)$ for a set with $\mathcal{I}$ real WD $i$ and a fixed set of allocation policy $\mathbb{R}$ indexed by $e\in\{1,\dots,\mathcal{E}\}$, our objective is to learn a permutation-equivariant mapping that outputs, for each $\mathbb{R}$, a probability distribution over $I$ plus a slack (unused capacity) entry, summing to one per $\mathbb{R}$. Formally,
@@ V-B Decoder
The SP1 decoder generates, for each resource $e$ in slice $n$, an augmented softmax distribution over the index set $\{0\}\cup\mathcal{I}$, where index 0 denotes a slack option corresponding to unused capacity. Let the resulting probabilities be $p_{0},p_{1},\ldots,p_{|\mathcal{I}|}$ with $p_{k}\geq 0$ and $\sum_{k}p_{k}=1$. Define $\phi^{n}_{i,e}=p_{i}$ for each user $i$. Then,
@@
Let the SP3 classifier output a probability vector $p_{i}=[p_{i}^{1},p_{i}^{2},\ldots,p_{i}^{K}]\in\mathbb{R}^{K}$, where $K=|\Delta_{i}|$ is the number of admissible offloading decisions for WD $i$. Since the final decision is taken as $\widehat{\delta}_{i}=\arg\max_{k}p_{i}^{k}$, exactly one index is active while all others are inactive. Therefore,
@@ VI-B Experiment Environment
We consider a small-scale scenario for an edge application with a square area of (500 x 500)$m$ where 8 COINs, 1 MEC, 3 APs, and WDs are placed at random. The power spectral density of Gaussian noise $N_{0}$ is 174dBm/Hz-174\ dBm/Hz [aliyu2024digital], and according to [jovsilo2022joint], we represent the achievable physical rate $R_{i,a}$ from WD $i$ to AP $a$ as $R_{i,a}=B_{a}log\ (1+(distance)P^{t}/N_{0})$, where distancedistance is the Euclidean distance between WD $i$ and AP $a$, and $P^{t}$ is the transmission power of WD $i$. Other simulation parameters are summarized in table II.
@@
  • Proportional Resource Allocation Policy (PRAP): We define the policies $\mathbb{R}_{\omega}^{pr}$, which share the bandwidth of each AP $a$ to each slice $n$ proportionally, $\mathbb{R}_{\phi_{a}}^{pr}$ which gives propotional share of the bandwidth of each AP $a$ in slice $n$ to each WD $i$, and finally $\mathbb{R}_{\phi_{j}}^{pr}$ which gives proportional share of the computing resource of node $j$ in slice $n$ to each WD $i$.
  • Equal Resource Allocation Policy (ERAP): We define the policies $\mathbb{R}_{\omega}^{eq}$, which gives equal share of the bandwidth of each AP $a$ to each slice $n$, $\mathbb{R}_{\phi_{a}}^{eq}$ which gives equal share of the bandwidth of each AP $a$ in slice $n$ to each WD $i$, and finally $\mathbb{R}_{\phi_{j}}^{eq}$ which gives equal share of the computing resource of node $j$ in slice $n$ to each WD $i$.
@@
To evaluate the overall efficiency of our proposed framework, we define the system cost ratio $CR$ as a key performance metric. The $CR$ quantifies the ratio between the system cost reached using the optimization solver ($C_{OS}$) and the system cost reached under the baseline model ($C_{m}$).
@@
A higher $CR$ indicates that the model is more effective at minimizing system costs and is performing closer to the optimal solution.
Figure 6 illustrates the $CR$ as a function of the number of APs for WD sizes of 15 and 25. In both scenarios, the proposed DeepSets-S model consistently achieves higher cost ratios than the baselines ERAP and PRAP, confirming its ability to approximate the optimal solution more closely. The performance gap is more pronounced for larger workloads (WD = 25), where DeepSets-S maintains a substantial advantage.